OpenStudy (anonymous):

check my answers to the following questions please simplify this expression (3/7)^2)My answer is 6/14 simplify this expression 3(14/(4^2-3^2)-5 = my answer 27 what is a solution to this equation y=2x+5) my answer (5,20)

OpenStudy (anonymous):

The first one is correct, but you can (and should) put it in simpler terms.

OpenStudy (campbell_st):

well if its \[(\frac{3}{7})^2 = \frac{3^2}{7^2}\] square the numerator and denominator squaring a number means multiplying the number by its self as an example \[5^2 = 5 \times 5 = 25\] hope it helps

OpenStudy (anonymous):

thankyou

OpenStudy (campbell_st):

the 2nd question seems to be incorrect as well its \[3 \times (\frac{14}{4^2 - 3^2}) - 5\] so start by simplifying the denominator of the fraction \[4^2 - 3^2 = ?\]

OpenStudy (anonymous):

4^2 =16 and 3^2=9